FIR SPECTROSCOPY OF BARE CH_{5}^{+} USING LASER INDUCED REACTIONS

\underline{OSKAR ASVANY}, EDOUARD HUGO, Leiden Observatory, P.O. Box 9513, 2300 RA Leiden, The Netherlands; STEPHAN SCHLEMMER, I. Physikalisches Institut, Universität zu Köln, Zülpicher Straße 77, 50937 Köln, Germany.

Laser Induced Reaction (LIR) is a novel technique for the spectroscopy of molecular ions as well as the study of statespecific rate coefficients and lifetimes of excited states. It is based on trapping mass-selected ions in the cold environment of a reactant gas and probing the laser-induced excitation of the ion by the outcome of the corresponding endothermic ion-molecule reaction. The feasibility of this method has been demonstrated in the past for the electronic spectroscopy of N_{2}^{+} in the charge transfer reaction with Ar, and the spectroscopy of the IR-active modes of the acetylene cation in the abstraction reaction of C_{2}H_{2}^{+} with hydrogen^{a}^{b}. Recently, LIR has been applied to record the first far infrared spectra of bare CH_{5}^{+} and some of its deuterated versions combining a low temperature 22–pole ion trap with the large tuning range of the free electron laser FELIX. To obtain the LIR spectrum, the product ions of the endothermic proton transfer reaction CH_{5}^{+} + CO_{2} ⟶ CH_{4} + OCOH^{+} have been counted as a function of the FELIX wavelength. In addition to the earlier observed C–H-stretching feature^{c} at 3000 cm^{−1}, low-frequency extensions down to 2200 cm^{−1} become discernible, as well as a H–C–H bending feature at 1250 cm^{−1}. Comparison of the measured data to computed ones shows that CH_{5}^{+} is fluxional at the temperature of the experiment, T =110K.

aS. Schlemmer, E. Lescop, J. v. Richthofen, and D. Gerlich, J. Chem. Phys. 117, 2068–2075 (2002)

bO. Asvany, T. Giesen, B. Redlich, and S. Schlemmer, Phys. Rev. Lett. 94, 073001, (2005)

cE. T. White, J. Tang, and T. Oka, Science 284, 135–137 (1999)